4. The market demand for a given commodity is denoted by QD = 1,000 - 10P. The market supply is captured by QS = 40P - 80 when P ≥ 2 and QS = 0 when P < 2. Answer the following questions.

a) Determine the equilibrium in this market and represent it on a properly labeled graph.

b) Represent on a graph the total surplus observed in this market and determine its dollar value. Is the resource allocation efficient in this case? Explain why or why not.

c) Suppose now that the market price is administratively set at P* = 25. Does this create excess demand or excess supply in the market? Quantify this shortage or surplus.

d) Represent on a graph the new total surplus observed in this market, as well as the deadweight loss experienced by society as a result of the pricing policy described in the previous question.
Answer #1

Qo= 1000-lop Qs = 40p as for equilibrum [Oo-os] 100o-lop240p 1000=sop IP = 420 → equilitim price Q = 40x 20 To = 800 equilibib) Total Surplus - Consumer & Surplus producer Surplus - Area of the triggle DEF = x base height = x 800 x 100 - Googleo ITSSurplus - Supply - Derrard (at p= 25) = 408-(1000 -lop - 40p flop-1000 = Sop-1000 - 50x25-1000 Suplus - 250 do Surplus is Cal